OpenStudy (anonymous):

Prove or Disprove: If m and n are distinct integers such that m,n >= 2 then the integers (mod m) and the integers (mod n) are disjoint.

OpenStudy (astrophysics):

@ganeshie8

OpenStudy (loser66):

To me, we disprove it, just take n = 2, m = 4, then set of integers =0 mod 4 and set of integers =0 mod 2 are joint. \(\{2k | k \in \mathbb Z\} \cap \{4l|l \in \mathbb Z\}\neq \emptyset \) Ex, \(4 \in \) both sets.

OpenStudy (anonymous):

The answer sheet tells me that this statement is true... I'm not sure why.

OpenStudy (anonymous):

Wouldn't the integers (mod m) mean { [0], [1], [2],...,[m-1]}.

zepdrix (zepdrix):

Yes, which is why I don't understand how it could possibly be true.\[\large\rm \mathbb Z_m=\{~0,1,2,3,...,m-1~\}\]\[\large\rm \mathbb Z_n=\{~0,1,2,3,...,n-1~\}\] So if you take like... m=4 and n=5,\[\large\rm \mathbb Z_4=\{~0,1,2,3~\}\]\[\large\rm \mathbb Z_5=\{~0,1,2,3,4~\}\]Clearly they share some stuff... I guess maybe I'm not understanding what they mean by disjoint. Does that refer to the elements? Err blah.. my math brain not working today :]

OpenStudy (anonymous):

I thought that the integers mod m meant the set of all the equivalence classes mod m and not the actual elements themselves.

zepdrix (zepdrix):

Hmm ya, you're probably right :) Sorry I'm a lil rusty. So then do we want to show that they share no equivalence class then? Is that what we're looking for with disjoint perhaps? Like for \(\large\rm \mathbb Z_5\) we would say that:\[\large\rm [1]=\{~1,6,11,...~\}\]But in \(\large\rm \mathbb Z_4\) we instead have:\[\large\rm [1]=\{~1,5,9,...~\}\]

OpenStudy (anonymous):

Something along those lines, I assume they want a way to show that no two equivalence classes are the same for all m and n.

OpenStudy (anonymous):

So for Z (mod 5) and Z (mod 4), they do not share any common elements such that there are no two equivalence classes that have the same elements in it.

OpenStudy (loser66):

Since m,n are arbitrary, if you take m = s*n, then pretty sure the equivalent classes are same, like this \([3] (mod 15) \equiv [1] (mod 3)\), clearly, 3 is not 15.